OpenStudy (anonymous):

Can somebody show how to integrate this? x^3(4-x^2)^1/2 dx

OpenStudy (irishboy123):

do it by parts ∫uv' = uv - ∫u'v v' = x(4-x^2)^1/2 so u = .....

OpenStudy (anonymous):

OMG!! You're so smart! Didnt think of that! Thanks!

OpenStudy (irishboy123):

it was just a suggestion. i hope it works ;)

OpenStudy (anonymous):

You can also use a trigonometric substitution. It's not pretty, but it does the job. Set \(x=2\sin u\), then \(dx=2\cos u\,du\), and \[\int x^3\sqrt{4-x^2}\,dx=\int 8\sin^3u\sqrt{4-4\sin^2u}\,(2\cos u)\,du=32\int \sin^3u\cos^2u\,du\] Now, \(\sin^3u\cos^2u=\sin u(\cos^2u-\cos^4u)\) by the Pythagorean identity, so the integral is equivalent to \[32\int \sin u(\cos^2u-\cos^4u)\,du\] which simplifies nicely with the substitution \(t=\cos u\) such that \(dt=-\sin u\,du\), giving \[32\int (t^4-t^2)\,dt\]